High school seniors' SAT scores are normally distributed with μ = 1050 and σ = 100

2.) High school seniors' SAT scores are normally distributed with μ = 1050 and σ = 100. If a student is selected at random, find the probability that her SAT score is: 

a.) above 1200 b.) below 890 c.) between 1000 and 1100 

d.) What SAT score separates the smartest 4% of students? 

e). If 18 seniors are selected, find the probability that their mean SAT score is above 1150 


3.) A survey of 200 college students revealed that 160 of them eat dessert daily. Construct a 95% confidence interval for the population proportion of college students who eat dessert daily
Answer #1

Hi, We are supposed to answer one question at a time. Since you have not mentioned which question to answer. I am answering the first one. Please repost the remaining question that you would like to be answered.

2.

Given Information:

Mean = 1050

Standard Deviation = 100

To calculate the probabilities, the z score needs to be calculated first using following formula:

X - u z= σ where, X= Target Value z=z score o = standard deviation

a) The probability that SAT score is above 1200

Z = 1200-1050 =1.5 100

A(z) probability of a normal random variable not being more than a standard deviations above its mean. Values of z of particu

The probability can be calculated as 1 - given area = 1 - 0.9332 = 0.0668

b) probability below 890

890-1050 = -1.6 II 100

.00 01 102 05 .09 -3.4 -3.3 -3.2 -3.1 -3.0 -2.9 -2.8 -2.7 -2.6 -2.5 -2.4 -2.3 -2.2 -2.1 -2.0 -1.9 -1.8 -1.7 -1.6 -1.5 -1.4 -1

In this case, the probability will be 0.0548

c) between 1000 to 1100

Z= 1000-1050 100 = -0.5

Z= 1100 - 1050 100 = 0.5

A(z) probability of a normal random variable not being more than a standard deviations above its mean. Values of z of particu

-1.2 -1.1 -1.0 -0.9 -0.8 -0.7 -0.6 0.5 -0.4 -0.3 -0.2 0.1 -0.0 1151 1357 1587 1.1841 12119 1.2420 1.2743 1.3085 3446 1.3821 .

The probability can be calculated as:

p = 0.6915 - 0.3085 = 0.383

d) The SAT score value that separates the 4% of students

The z score at p - 0.04 is 1.75

X – 1050 1.75 = 100 X = 1050 +1.75*100 = 1225

Therefore, the SAT score is 1225

e) following is the formula for calculating z score when sample size is given

X – u Z= o Jn where, X= Target Value z=z score o = standard deviation = sample size

Z= = = 4.24 1150-1050 100 18

By referring to the standard normal table, the p-value at z = 4.24 is 0.0001